OpenStudy (kaylala):

Find the value of the following using the concepts of both addition and subtraction formulas and reference angle: tan (-165)

OpenStudy (anonymous):

tan(-165) = - tan(165) = - tan(180-165) = - ( -tan(15)) = tan 15

OpenStudy (kaylala):

okay

OpenStudy (kaylala):

= 2-(sqrt3) ? @MonoPolyRoly

OpenStudy (kaylala):

did i get it right?

OpenStudy (anonymous):

Yes.

OpenStudy (kaylala):

okay thanks @MonoPolyRoly :)

OpenStudy (agent0smith):

you could also just find -tan(120+45) then use the angle sum formula

OpenStudy (kaylala):

what's that? i dont understand. really sorry @agent0smith

OpenStudy (agent0smith):

what's what...? " using the concepts of both addition and subtraction formulas" you're supposed to use angle sum formulas

OpenStudy (anonymous):

Hmm, when he finds tan(15), can't he use tan(45-30) ?

OpenStudy (kaylala):

oh i see now @agent0smith but was the answer that i found correct?

OpenStudy (agent0smith):

Yeah, i'm just saying you can skip a couple of steps. Same thing either way really... and it does say to use a reference angle.

OpenStudy (anonymous):

You can have many ways to do this question...

OpenStudy (kaylala):

could you please explain: - ( -tan(15)) i dont understand how this happened? @MonoPolyRoly

OpenStudy (anonymous):

tan(-165) = -tan(165) = -tan(120+45) Use compound angle formula. Note that tan(120) = - tan(180- 60) = - ( - tan(60) ). Here, you use reference angle.

OpenStudy (anonymous):

Sorry, my mistake in my first comment here. It should be tan(-165) = - tan(165) = - tan(180-15) = - ( -tan(15)) = tan 15 Note that tan(180-x) = -(tanx).

OpenStudy (anonymous):

You may find this table helpful.

OpenStudy (kaylala):

what happens now? will the answer change @MonoPolyRoly

OpenStudy (anonymous):

The answer won't change. It's just that I made a silly mistake in the part " - tan(165) = - tan(180-15)". I wrote " - tan(165) = - tan(180-165)", which was not correct, as tan(165) ≠ tan(180-165).

OpenStudy (kaylala):

oh okay.